Marty pays each of his workers $0.80 per day. The cost of his other variable inputs is $0.50 per cup of yogurt. His fixed cost is $100 per day.

What are Marty's variable cost and total cost when he produces 110 cups of yogurt? 200 cups?
1
Expert's answer
2012-03-15T09:57:02-0400
If Marty produces 110 cups of yougurt per day:
Martsy`s other& variable costs $0.50*110=$55.0
If Marty produces 200 cups of yougurt per day:
Martsy`s other& variable costs $0.50*200=$100.0

Fixed cost per day = $100
In case Marty has one worker, than cost for workers will be $0.80 per day
In case Marty has more than one worker than cost will be $0.80*N, N-quantity of workers

FOR 110 cups:
Variable cost =$55.0 +$0.80*N
TOTAL Cost = $55.0+$0.80*N + $100= $155+$0.80*N

FOR 200 cups:
Variable cost =$100.0 +$0.80*N
TOTAL Cost = $100.0+$0.80*N + $100= $200+$0.80*N